# Statistical Inference for Explainable Boosting Machines

### v1

@@ Appendix A Additional Literature
(Wahba, 1983) established the asymptotic distribution of smoothing spline estimators via a Bayesian posterior covariance argument, (Silverman, 1985) developed the equivalent kernel representation and convergence theory, and (Nychka, 1988) analyzed the frequentist coverage of Bayesian confidence intervals for smoothing splines. Built on these developments, (Wood, 2006) investigated the coverage of penalized regression spline intervals in the GAM setting, while (Yoshida and Naito, 2012) provided a full asymptotic distribution theory for penalized spline estimators, establishing asymptotic normality for each additive component.
Appendix B Algorithms
@@
1:Input: $\mathbf{X}\in\mathbb{R}^{n\times p}$, $\mathbf{y}\in\mathbb{R}^{n}$, subsample rate $\xi$, rounds $T$, truncation level $M>0$.
@@
1:Input: $\mathbf{X}\in\mathbb{R}^{n\times p}$, $\mathbf{y}\in\mathbb{R}^{n}$, subsample rate $\xi$, rounds $T$, truncation level $M>0$.
@@
Appendix C Binning
C.1 Bin-level Decomposition of the Structure Matrix
@@
Lemma C.1 (Bin-Level Decomposition of Structure Matrix).
@@
C.2 Computing the Kernel Vector in Bin Space
@@
$\mathbf{k}^{(k)}_{n}(\mathbf{x})^{:}=\mathbb{E}\left[\mathbf{b}^{(k)}(\mathbf{x})^{\top}\mathbf{D}^{(k)}{\mathbf{b}_{n}^{(k)}}^{\top}\right]\in\mathbb{R}^{1\times n},\quad\mathbf{K}^{(k)}_{n}:=\mathbb{E}[\mathbf{S}^{(k)}_{n}]=\mathbb{E}\left[\mathbf{b}_{n}^{(k)}\mathbf{B}^{(k)}{\mathbf{b}_{n}^{(k)}}^{\top}\right],\quad\mathcal{K}=\sum_{k=1}^{p}\mathbb{E}\left[\mathbf{b}^{(k)}_{n}\mathbf{B}^{(k)}{\mathbf{b}^{(k)}_{n}}^{\top}\right$],
@@
$\mathbf{r}^{(k)}(\mathbf{x})^{\top}:=\mathbb{E}[{\mathbf{b}_{n}^{(k)}}\mathbf{D}^{(k)}\mathbf{b}^{(k)}(\mathbf{x})]^{\top}\mathbf{J}_{n}[\mathbf{I}+\mathbf{J}_{n}\mathcal{K}]^{-1},$
@@
Lemma C.2 (Norm of the weight vector in bin-space).
@@
$\mathcal{K}=\sum_{k=1}^{p}\mathbb{E}\left[\mathbf{b}^{(k)}_{n}\mathbf{B}^{(k)}{\mathbf{b}^{(k)}_{n}}^{\top}\right]=\mathbf{Z}_{n}^{(k)}\left(\sum_{k=1}^{p}\mathbb{E}\left[\mathbf{D}^{(k)}\mathbf{B}^{(k)}\mathbf{D}^{(k)}\right]\right){\mathbf{Z}_{n}^{(k)}}^{\top}$
@@
$\mathbf{r}^{(k)}_{n}(\mathbf{x})^{\top}$ $:=\mathbb{E}[{\mathbf{b}_{n}^{(k)}}\mathbf{D}^{(k)}\mathbf{b}^{(k)}(\mathbf{x})]^{\top}\mathbf{J}_{n}[\mathbf{I}+\mathbf{J}_{n}\mathcal{K}]^{-1}$
@@
Lemma C.3 (Bin-space norm computation).
@@
Appendix D Finite Sample Convergence
D.1 Structural assumptions and identities for finite-sample convergence
@@
Lemma D.1 (Centered+ones decomposition and identity on $\mathcal{S}$).
@@
Lemma D.1 supplies the identity-on-$\mathcal{S}$ bound
@@
D.2 A Fixed-point Iteration
@@
D.3 Another Fixed-point Iteration
@@ D.4.1 Algorithm 1
D.4.2 Algorithm 2
Under stated assumptions, we can prove that Algorithm 1’s prediction converge almost surely to a postulated fixed point. Formally, we make use of Theorem F.1, called a stochastic contraction mapping theorem, to demonstrate the difference between the current prediction and the fixed point is a stochastic contraction mapping.
Theorem D.4.
@@ D.4.2 Algorithm 2
Appendix E Limiting Distribution
In Section D, we have shown that as we keep boosting, our feature estimators will converge to a stacked kernel-ridge-regression
@@ Appendix E Limiting Distribution
E.1 Preliminaries for Asymptotic Normality
@@
The first term can be easily shown by constraining leaf size, which we discuss in Section E.2. The later term is trickier. Generalizing from training set to the whole input space, we expect a convergence that is at least of a weak. Bounded kernel, fortunately, allows us to do so.
E.2 Conditional Asymptotic Normality
@@
E.2.1 Subsampling
The key component of this proof is the rate of $\left\|\mathbf{r}^{(k)}_{E}\right\|_{1},\left\|\mathbf{r}^{(k)}_{E}\right\|,\left\|\mathbf{r}^{(k)}_{E}\right\|_{\infty}$. These rates will be built on a result shown in (Zhou and Hooker, 2019) on the behavior of subsampling rate.
Lemma E.1.
@@ E.2.1 Subsampling
For each algorithm, our plan to conditional CLT is as follow: Lemma E.1 specifies a proper subsample rates that enables norm convergence. Lindeberg condition simply asks us to check the truncation level $\frac{\mathbf{r}^{(k)}_{i}}{\left\|\mathbf{r}^{(k)}\right\|}$. We will first study the norm of $\mathbf{k}^{(k)}$. To bound the $\ell_{1}$ norm, it suffices to imposing conditions on the leaf size (with respect to the Lebesgue measure of data points). To bound the $\ell_{2}$ norm. A simple algebra utilizing the rate of $\ell_{1}$ norm will help with that, too.If this norm is well controlled, notice that in both algorithm, the Boulevard vector $\mathbf{r}^{(k)}$ is simply a sum of rows in the matrix $\mathbf{P}_{1}:=\mathbf{J}_{n}[\mathbf{I}+\mathbf{J}_{n}\mathbf{K}]^{-1}$, $\mathbf{P}_{2}:=(\mathbf{I}-\mathbf{K}^{(k)})^{-1}(\mathbf{I}-\frac{1}{n}\mathbf{1}\mathbf{1}^{\top})\Big[\mathbf{I}-(\mathbf{I}+\mathcal{K})^{\dagger}\mathcal{K}\Big]$ with the weighting vector being $\mathbf{k}^{(k)}$. Hence it suffices to study the eigenvalues of the operator $\mathbf{P}_{1},\mathbf{P}_{2}$, which we discuss below. We will first show that the spectrum of $\mathbf{K}^{(k)}$ is strictly bounded way from 1.
E.2.2 Expected Feature Tree Kernels
Lemma E.2 (Spectrum of $\mathbf{K}^{(k)}$).
@@ E.2.2 Expected Feature Tree Kernels
Lemma E.2 gives a strict bound on the eigenvalues of the expected kernel, which we can utilize to stabilize the behavior of the aggregated kernel $\mathbf{K}$ and $\mathcal{K}$, which is described in the following sections.
E.2.3 Operator $\mathbf{P}_{1}$ for Algorithm 1
@@
Corollary E.3 (Eigenvalue bounds for $\mathbf{P}_{1}$).
@@
E.2.4 Operator $\mathbf{P}_{2}$ for Algorithm 2
Lemma E.4.
@@ E.2.4 Operator $\mathbf{P}_{2}$ for Algorithm 2
Lemma E.5 (Eigenvalues of $\mathbf{P}_{2}$).
@@ E.2.5 Rate of Convergence
Lemma E.6 (Rate of Convergence).
@@
implies that $\mathbf{k}^{(k)}_{nj}=O(n^{-2/3})$. By Lemma E.1, $\left\|\mathbf{k}^{(k)}\right\|_{1}\leq 1$, then an $\ell_{2}$ norm bound yields
@@
By corollary E.3 and lemma E.5, we know that $\mathbf{r}^{(k)}_{E}$ are linear transformations from $\mathbf{k}^{(k)}$ by $\mathbf{P}_{1},\mathbf{P}_{2}$ with bounded eigenvalues. Hence $\left\|\mathbf{r}^{(k)}_{E}\right\|$ enjoys similar rates. ∎
@@
Theorem E.7 (Conditional Asymptotic Normality).
@@
E.3 Bias Control and Asymptotic Unbiasedness
In Section E.2 we established a conditional central limit theorem: conditioning on the training covariates $\mathbf{X}$, the noise contribution $\mathbf{r}^{(k)}(\mathbf{x})^{\top}\vec{\epsilon}$ is asymptotically normal with variance $\sigma^{2}$. To extend this to an unconditional CLT for the estimator, it remains to verify that all bias terms arising from the random design vanish at rate $o_{p}(1)$.
@@ E.3 Bias Control and Asymptotic Unbiasedness
E.3.1 Baseline Bias
Lemma E.8 (Baseline Bias).
Let $\beta\in\mathbb{R}^{n}$ denote the baseline vector. With subsampling rate specified in Section E.2.1, $\mathbf{r}^{(k)}_{E}(\mathbf{x})^{\top}\beta=O(\frac{1}{n})$ for all $k=1,\dots,p$.
@@ E.3.1 Baseline Bias
The same feature term bias can be killed by building deeper trees as we collect more and more observations from the underlying data generating process. This can be realized by inheriting assumptions on leaf size from (Zhou and Hooker, 2019), which we present below:
E.3.2 Same-feature Bias for Algorithm 1
Lemma E.9 (Same–feature Bias for Algorithm 1).

### v2

@@ Appendix B Additional Literature
Wahba (1983) established the asymptotic distribution of smoothing spline estimators via a Bayesian posterior covariance argument, Silverman (1985) developed the equivalent kernel representation and convergence theory, and Nychka (1988) analyzed the frequentist coverage of Bayesian confidence intervals for smoothing splines. Built on these developments, Wood (2006) investigated the coverage of penalized regression spline intervals in the GAM setting, while Yoshida and Naito (2012) provided a full asymptotic distribution theory for penalized spline estimators, establishing asymptotic normality for each additive component.
Appendix C Algorithms
@@
1:Input: $\mathbf{X}\in\mathbb{R}^{n\times p}$, $\mathbf{y}\in\mathbb{R}^{n}$, subsample rate $\xi$, rounds $B$, truncation level $M>0$.
@@
1:Input: $\mathbf{X}\in\mathbb{R}^{n\times p}$, $\mathbf{y}\in\mathbb{R}^{n}$, subsample rate $\xi$, rounds $B$, truncation level $M>0$.
@@
Appendix D Binning
D.1 Bin-level Decomposition of the Structure Matrix
@@
Lemma D.1 (Bin-Level Decomposition of Structure Matrix).
@@
D.2 Computing the Kernel Vector in Bin Space
@@
$\mathbf{k}^{(k)}_{n}(\mathbf{x})^{:}=\mathbb{E}\left[\mathbf{b}^{(k)}(\mathbf{x})^{\top}\mathbf{D}^{(k)}{\mathbf{b}_{n}^{(k)}}^{\top}\right]\in\mathbb{R}^{1\times n},\quad\mathbf{K}^{(k)}_{n}:=\mathbb{E}[\mathbf{S}^{(k)}_{n}]=\mathbb{E}\left[\mathbf{b}_{n}^{(k)}\mathbf{B}^{(k)}{\mathbf{b}_{n}^{(k)}}^{\top}\right],\quad\mathcal{K}_{\text{bin}}=\sum_{k=1}^{p}\mathbb{E}\left[\mathbf{b}^{(k)}_{n}\mathbf{B}^{(k)}{\mathbf{b}^{(k)}_{n}}^{\top}\right$],
@@
$\mathbf{r}^{(k)}(\mathbf{x})^{\top}:=\mathbb{E}[{\mathbf{b}_{n}^{(k)}}\mathbf{D}^{(k)}\mathbf{b}^{(k)}(\mathbf{x})]^{\top}\mathbf{J}_{n}[\mathbf{I}+\mathbf{J}_{n}\mathcal{K}_{\text{bin}}]^{-1},$
@@
Lemma D.2 (Norm of the weight vector in bin-space).
@@
$\mathcal{K}_{\text{bin}}=\sum_{k=1}^{p}\mathbb{E}\left[\mathbf{b}^{(k)}_{n}\mathbf{B}^{(k)}{\mathbf{b}^{(k)}_{n}}^{\top}\right]=\mathbf{Z}_{n}^{(k)}\left(\sum_{k=1}^{p}\mathbb{E}\left[\mathbf{D}^{(k)}\mathbf{B}^{(k)}\mathbf{D}^{(k)}\right]\right){\mathbf{Z}_{n}^{(k)}}^{\top}$
@@
$\mathbf{r}^{(k)}_{n}(\mathbf{x})^{\top}$ $:=\mathbb{E}[{\mathbf{b}_{n}^{(k)}}\mathbf{D}^{(k)}\mathbf{b}^{(k)}(\mathbf{x})]^{\top}\mathbf{J}_{n}[\mathbf{I}+\mathbf{J}_{n}\mathcal{K}_{\text{bin}}]^{-1}$
@@
Lemma D.3 (Bin-space norm computation).
@@
D.3 What we lose by binning.
Binning induces a discretization: all points within the same histogram bin are treated as indistinguishable along that feature. This can (i) introduce an approximation bias at sub-bin scales, and (ii) prevent recovery of within-bin variation even when $n$ is large. However, our bin-space computations do not introduce an additional approximation beyond the estimator we analyze. The reasoning is as follow: the practical EBM implementation we build on already trains histogram trees, so the learned model is itself defined on bins, and our inference targets the sampling variability of this histogram-EBM estimator.
If one instead wants inference for a hypothetical unbinned (continuous) version of the model, an additional discretization error must be accounted for. A simple bound can be stated under the smoothness assumption in Assumption 2.1. Suppose the bin width along each feature is $O(1/m)$. Then an application of the mean value inequality yields a uniform approximation error $\|\hat{f}-\hat{f}^{(\mathrm{bin})}\|_{\infty}=O(1/m)$ (assuming the relevant derivatives/densities are bounded), so the accumulated increase in MSE scales as $O(p/m)$. In particular, as the bin resolution increases (e.g., $m\to\infty$ with $m\gg n^{1/3}$), this discretization effect becomes negligible, recovering the same MSE order as in the unbinned idealization.
Appendix E Finite Sample Convergence
E.1 Structural assumptions and identities for finite-sample convergence
@@
Lemma E.1 (Centered+ones decomposition and identity on $\mathcal{S}$).
@@
Lemma E.1 supplies the identity-on-$\mathcal{S}$ bound
@@
E.2 A Fixed-point Iteration
@@
E.3 Another Fixed-point Iteration
@@ E.4.1 Algorithm 1
E.4.2 Algorithm 2
Under stated assumptions, we can prove that Algorithm 1’s prediction converge almost surely to a postulated fixed point. Formally, we make use of Theorem G.1, called a stochastic contraction mapping theorem, to demonstrate the difference between the current prediction and the fixed point is a stochastic contraction mapping.
Theorem E.4.
@@ E.4.2 Algorithm 2
Empirical verification of Assumption 4.9 (Partition Distinctness).
In the proof above we use Assumption 4.9 to control the bound of the operator norm of the cross-feature blocks $P_{k}=\mathbb{E}[\mathbf{J}_{n}\mathbf{S}^{(k)}\mathbf{J}_{n}]$, i.e. $c_{3}=\max_{a\neq k}\|P_{a}P_{k}\|$ and ensure $(p-1)c_{3}<1$. Table 10 reports empirical estimates of $c_{3}$ and the corresponding $(p-1)c_{3}$ across representative real and synthetic datasets.
Table 10: Empirical verification of Assumption 4.9 across real and synthetic datasets. All experiments use learning_rate=0.1, max_rounds=200, subsample_rate=0.7, truncation=3.0.
Dataset n p $\mathbf{c_{3}}$ $\mathbf{(p-1)c_{3}}$
adult 48842 14 $5.227\times 10^{-4}$ $6.795\times 10^{-3}$
bike_sharing 17379 13 $9.513\times 10^{-5}$ $1.142\times 10^{-3}$
energy_efficiency 768 8 $2.838\times 10^{-4}$ $1.987\times 10^{-3}$
breast_cancer 569 30 $5.553\times 10^{-2}$ $1.610\times 10^{+0}$
california_housing 20640 8 $4.434\times 10^{-5}$ $3.104\times 10^{-4}$
$\mathbf{y}=0.9\mathbf{X}+\sqrt{1-0.81}\epsilon,\epsilon\sim\mathcal{N}(0,1)$ 5000 10 $7.181\times 10^{-4}$ $6.463\times 10^{-3}$
Across datasets we observe that the cross-products are typically negligible. There are several potential theories for such observations:
  1. 1.
    While multicollinearity is almost surely present in most datasets, their statistical dependence can be weak across features after centering and preprocessing. This is widely noted in GAM literature: Hastie and Tibshirani (1986) often assumes low dependence so that model specifications do not run into identifiability issues.
  2. 2.
    Even when features are correlated, the resulting feature-wise kernels can still point to different directions in the RKHS, which results in a zero cross-product in expectation. Here is a simple GAM as an example: $\mathbf{y}=\mathbf{f}_{1}(\mathbf{X}_{1})+\mathbf{f}_{2}(\mathbf{X}_{2})+\epsilon$, where $\mathbf{f}_{1}(\mathbf{x})=\mathbf{x},\mathbf{f}_{2}(\mathbf{x})=\mathbf{x}^{2}+1$, and $(\mathbf{X}_{1},\mathbf{X}_{2})$ are jointly Gaussian with non-zero correlation. The covariance $\text{Cov}\left(\mathbf{f}_{1}(\mathbf{X}_{1}),\mathbf{f}_{2}(\mathbf{X}_{2})\right)$ can be shown to equal to 0 for any $\rho$. So unless two covariates are nearly identical in both their distribution and main effects to the signal, the induced smoothing kernel can theoretically be small.
  3. 3.
    Last but not least, Gradient Boosting learns by approximating residuals, which are approximately orthogonal to the previously fitted learners.
Appendix F Limiting Distribution
In Section E, we have shown that as we keep boosting, our feature estimators will converge to a stacked kernel-ridge-regression
@@ Appendix F Limiting Distribution
F.1 Preliminaries for Asymptotic Normality
@@
The first term can be easily shown by constraining leaf size, which we discuss in Section F.2. The later term is trickier. Generalizing from training set to the whole input space, we expect a convergence that is at least of a weak. Bounded kernel, fortunately, allows us to do so.
F.2 Conditional Asymptotic Normality
@@
F.2.1 Subsampling
The key component of this proof is the rate of $\left\|\mathbf{r}^{(k)}_{E}\right\|_{1},\left\|\mathbf{r}^{(k)}_{E}\right\|,\left\|\mathbf{r}^{(k)}_{E}\right\|_{\infty}$. These rates will be built on a result shown in Zhou and Hooker (2019) on the behavior of subsampling rate.
Lemma F.1.
@@ F.2.1 Subsampling
For each algorithm, our plan to conditional CLT is as follow: Lemma F.1 specifies a proper subsample rates that enables norm convergence. Lindeberg condition simply asks us to check the truncation level $\frac{\mathbf{r}^{(k)}_{i}}{\left\|\mathbf{r}^{(k)}\right\|}$. We will first study the norm of $\mathbf{k}^{(k)}$. To bound the $\ell_{1}$ norm, it suffices to imposing conditions on the leaf size (with respect to the Lebesgue measure of data points). To bound the $\ell_{2}$ norm. A simple algebra utilizing the rate of $\ell_{1}$ norm will help with that, too.If this norm is well controlled, notice that in both algorithm, the Boulevard vector $\mathbf{r}^{(k)}$ is simply a sum of rows in the matrix $\mathbf{P}_{1}:=\mathbf{J}_{n}[\mathbf{I}+\mathbf{J}_{n}\mathbf{K}]^{-1}$, $\mathbf{P}_{2}:=(\mathbf{I}-\mathbf{K}^{(k)})^{-1}(\mathbf{I}-\frac{1}{n}\mathbf{1}\mathbf{1}^{\top})\Big[\mathbf{I}-(\mathbf{I}+\mathcal{K})^{\dagger}\mathcal{K}\Big]$ with the weighting vector being $\mathbf{k}^{(k)}$. Hence it suffices to study the eigenvalues of the operator $\mathbf{P}_{1},\mathbf{P}_{2}$, which we discuss below. We will first show that the spectrum of $\mathbf{K}^{(k)}$ is strictly bounded way from 1.
F.2.2 Expected Feature Tree Kernels
Lemma F.2 (Spectrum of $\mathbf{K}^{(k)}$).
@@ F.2.2 Expected Feature Tree Kernels
Lemma F.2 gives a strict bound on the eigenvalues of the expected kernel, which we can utilize to stabilize the behavior of the aggregated kernel $\mathbf{K}$ and $\mathcal{K}$, which is described in the following sections.
F.2.3 Operator $\mathbf{P}_{1}$ for Algorithm 1
@@
Corollary F.3 (Eigenvalue bounds for $\mathbf{P}_{1}$).
@@
F.2.4 Operator $\mathbf{P}_{2}$ for Algorithm 2
Lemma F.4.
@@ F.2.4 Operator $\mathbf{P}_{2}$ for Algorithm 2
Lemma F.5 (Eigenvalues of $\mathbf{P}_{2}$).
@@ F.2.5 Rate of Convergence
Lemma F.6 (Rate of Convergence).
@@
implies that $\mathbf{k}^{(k)}_{nj}=O(n^{-2/3})$. By Lemma F.1, $\left\|\mathbf{k}^{(k)}\right\|_{1}\leq 1$, then an $\ell_{2}$ norm bound yields
@@
By corollary F.3 and lemma F.5, we know that $\mathbf{r}^{(k)}_{E}$ are linear transformations from $\mathbf{k}^{(k)}$ by $\mathbf{P}_{1},\mathbf{P}_{2}$ with bounded eigenvalues. Hence $\left\|\mathbf{r}^{(k)}_{E}\right\|$ enjoys similar rates. ∎
@@
Theorem F.7 (Conditional Asymptotic Normality).
@@
F.3 Bias Control and Asymptotic Unbiasedness
In Section F.2 we established a conditional central limit theorem: conditioning on the training covariates $\mathbf{X}$, the noise contribution $\mathbf{r}^{(k)}(\mathbf{x})^{\top}\vec{\epsilon}$ is asymptotically normal with variance $\sigma^{2}$. To extend this to an unconditional CLT for the estimator, it remains to verify that all bias terms arising from the random design vanish at rate $o_{p}(1)$.
@@ F.3 Bias Control and Asymptotic Unbiasedness
F.3.1 Baseline Bias
Lemma F.8 (Baseline Bias).
Let $\beta\in\mathbb{R}^{n}$ denote the baseline vector. With subsampling rate specified in Section F.2.1, $\mathbf{r}^{(k)}_{E}(\mathbf{x})^{\top}\beta=O(\frac{1}{n})$ for all $k=1,\dots,p$.
@@ F.3.1 Baseline Bias
The same feature term bias can be killed by building deeper trees as we collect more and more observations from the underlying data generating process. This can be realized by inheriting assumptions on leaf size from Zhou and Hooker (2019), which we present below:
F.3.2 Same-feature Bias for Algorithm 1
Lemma F.9 (Same–feature Bias for Algorithm 1).